OpenStudy (anonymous):

This is confusing me 5(c-4)-2c=1

OpenStudy (luigi0210):

Distribute the 5

OpenStudy (anonymous):

5c-20-2c=1

OpenStudy (luigi0210):

Yea, now simplify

OpenStudy (anonymous):

but when i got to here i got confused like is it 7c-20=1?

OpenStudy (luigi0210):

It's 5c-2c, not 5c+2c

OpenStudy (anonymous):

Ohhhh so its 3c-20=1

OpenStudy (skullpatrol):

Yes.

OpenStudy (luigi0210):

Do you know what to do now?

OpenStudy (anonymous):

omgg ty so much :)

OpenStudy (anonymous):

Yes ty ^^

OpenStudy (luigi0210):

Alright, good luck!

OpenStudy (skullpatrol):

What answer did you get?

OpenStudy (anonymous):

c=7 :)

OpenStudy (skullpatrol):

Good job :D

OpenStudy (luigi0210):

Good job :P

OpenStudy (anonymous):

Ty i feel so smart now XD

OpenStudy (skullpatrol):

Remember to check your answer in the original equation...

OpenStudy (anonymous):

ok :3

OpenStudy (anonymous):

umm i need help with something else :c

OpenStudy (skullpatrol):

quick?

OpenStudy (anonymous):

it says Three less than the quotient of a number and 5 i write it as 3-x/5

OpenStudy (skullpatrol):

Be very careful with the phrase "less than" when talking about subtraction, read it as "Three less, than " so the words can not be translated in order.

OpenStudy (anonymous):

ok

OpenStudy (skullpatrol):

$$\huge \frac{x}{5}-3$$

OpenStudy (anonymous):

ohhh

OpenStudy (skullpatrol):

3 less, than a quantity is Q - 3

OpenStudy (anonymous):

k